### v1

# Spontaneous Magnon Decays from Nonrelativistic Time-Reversal Symmetry Breaking in Altermagnets

Spontaneous Magnon Decays from Nonrelativistic Time-Reversal Symmetry Breaking in Altermagnets
Rintaro Eto Department of Applied Physics, Waseda University, Okubo, Shinjuku-ku, Tokyo 169-8555, Japan Institut für Physik, Johannes Gutenberg Universität Mainz, D-55099 Mainz, Germany    Matthias Gohlke Theory of Quantum Matter Unit, Okinawa Institute of Science and Technology Graduate University, Onna-son, Okinawa 904-0495, Japan Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Str. 38, 01187 Dresden, Germany    Jairo Sinova Institut für Physik, Johannes Gutenberg Universität Mainz, D-55099 Mainz, Germany    Masahito Mochizuki Department of Applied Physics, Waseda University, Okubo, Shinjuku-ku, Tokyo 169-8555, Japan    Alexander L. Chernyshev Department of Physics and Astronomy, University of California, Irvine, California 92697, USA    Alexander Mook Institut für Physik, Johannes Gutenberg Universität Mainz, D-55099 Mainz, Germany
(February 27, 2025)
Quasiparticles are central to condensed matter physics, but their stability can be undermined by quantum many-body interactions. Magnons, quasiparticles in quantum magnets, are particularly intriguing because their properties are governed by both real and spin space. While crystal symmetries may be low, spin interactions often remain approximately isotropic, limiting spontaneous magnon decay. Textbook wisdom holds that collinear Heisenberg magnets follow a dichotomy: ferromagnets host stable magnons, while antiferromagnetic magnons may decay depending on dispersion curvature. Up to now, relativistic spin-orbit coupling and noncollinear order that connect spin space to real space, were shown to introduce more complex magnon instability mechanisms. Here, we show that even in nonrelativistic isotropic collinear systems, this conventional dichotomy is disrupted in altermagnets. Altermagnets, a newly identified class of collinear magnets, exhibit compensated spin order with nonrelativistic time-reversal symmetry breaking and even-parity band splitting. Using kinematic analysis, nonlinear spin-wave theory, and quantum simulations, we reveal that even weak band splitting opens a decay phase space, driving quasiparticle breakdown. Additionally, d-wave altermagnets form a rare “island of stability” at the Brillouin zone center. Our findings establish a quasiparticle stability trichotomy in collinear Heisenberg magnets and position altermagnets as a promising platform for unconventional spin dynamics.
The concept of a quasiparticle—an emergent, long-lived excitation that weakly interacts with its environment—is foundational to condensed matter physics, shaping our understanding of diverse systems [1, 2]. A central question is when and how quasiparticles break down at zero temperature due to quantum fluctuations [3, 4]. In bosonic systems, many-body interactions can trigger spontaneous quasiparticle decay, as seen in the instability of Bogoliubov excitations [5, 6] and anharmonic phonon decay in solids [7]. In quantum magnets, magnons—collective spin-wave excitations—are subject to unavoidable nonlinear interactions, leading to pronounced decay in frustrated and low-dimensional systems [3]. Understanding these decay mechanisms is crucial for determining whether quasiparticles remain the appropriate low-energy degrees of freedom or give way to more complex excitations. Experimentally, such decay manifests as spectral broadening and weight redistribution in neutron scattering and optical spectroscopy, offering insight into a material’s quantum many-body physics.
In many magnetically ordered insulators, spin space remains approximately isotropic and their leading-order physics is described by the paradigmatic Heisenberg model. This is the case when relativistic effects are negligible, such that spin space and real space are not connected. Once the spontaneous magnetic long-range order breaks continuous spin-rotation symmetry, gapless Nambu-Goldstone modes and universal low-energy behavior emerge. In collinear magnets—the simplest form of magnetic order—a dichotomy of magnon stability emerges: ferromagnetic magnons are inherently stable, as their ground state and excitations are exact eigenstates free of quantum fluctuations, whereas antiferromagnetic magnons experience fluctuations that can induce spontaneous decay [8, 3]. This dichotomy collapses when spin space becomes connected to real space by the relativistic spin-orbit coupling, noncollinear order, or dipolar interactions, making magnon decay a more intricate problem [9, 10, 11, 3, 12, 13, 14, 15, 16, 17, 18].
Altermagnets—a newly identified class of collinear magnets [19, 20]—challenge this dichotomy even within the nonrelativistic limit. Like antiferromagnets, they feature compensated magnetic order and quantum fluctuations, yet they also break time-reversal symmetry and exhibit nonrelativistic band splitting reminiscent of ferromagnets. Their unconventional beyond-s-wave even-parity spin splitting (d-, g-, or i-wave) of magnons [21, 22] raises a fundamental question: Can altermagnets support stable magnons, and how does their distinctive spin splitting interact with quantum many-body fluctuations to influence magnon stability? This question has gained urgency following quantum simulations of spin-$1/2$ altermagnets [23], which suggest pronounced nonclassical effects and high-energy magnon instability.
Figure 1: Stability of low-energy magnons in the three classes of collinear magnets in the nonrelativistic limit of an isotropic Heisenberg model. Due to spin conservation, magnons carry a spin quantum number ±$\pm 1$, indicated by color, and potential decay processes have to conserve spin. (a) Class-(i) magnets: Magnets with s-wave spin splitting, such as ferromagnets and ferrimagnets, exhibit inherently stable magnons at low energies. In ferromagnets, there are no quantum fluctuations and there is only one (acoustic) magnon branch, which can be considered the limit of infinite s-wave splitting. In contrast, ferrimagnets fluctuate. They support two oppositely polarized sublattices with unequal moments, and their spin splitting, i.e., the gap $\Delta_{0}$, is finite. It prohibits magnon decay at energies below $\Delta_{0}$, ensuring their stability. The indicated decay process of the magenta magnon is only possible for high-energy magnons above $\Delta_{0}$, such that one decay product can carry opposite spin (cyan). Thus, the lower magnon branch is stable below the gap $\Delta_{0}$, but the upper magnon branch can decay. (b) Class-(ii) magnets: Antiferromagnets fluctuate and have spin-degenerate magnons. Stability depends on the curvature of the magnon spectrum: negatively curved spectra ($\kappa<0$) lead to stable magnons, while positively curved spectra ($\kappa>0$) result in instability. (c) Class-(iii) magnets: Altermagnets are characterized by unconventional time-reversal symmetry breaking and beyond-s-wave even-parity spin splitting such as d/g/i-wave. In nodal planes, i.e., directions without spin splitting, the stability mirrors that of class-(ii) magnets, where the curvature of the spectrum dictates whether magnons can decay, but is additionally modified by the altermagnetic splitting, as indicated by the asterisk. Along directions with finite spin splitting the upper magnon branch is universally unstable due to the indicated decay process. The lower magnon branch exhibits rich stability physics that depends on the symmetry of the splitting as indicated by the asterisk: In d-wave magnets, it is either completely stable for $\kappa<0$ or exhibits a d-wave-shaped island of stability in the vicinity of the origin for $\kappa>0$. In contrast, there is no such island of stability for general g-wave and i-wave altermagnets.
Here, we combine kinematic analysis, many-body perturbation theory within nonlinear spin-wave theory, and nonperturbative quantum simulations to demonstrate that magnons in altermagnets are generally unstable, with exceptions to be discussed below. Even minimal even-parity spin splitting generically opens a finite phase space for decay, rendering the high-energy magnon branch universally unstable. Strikingly, the symmetry of the spin splitting is crucial for the stability of the lower-energy magnon branch: while d-wave spin splitting can protect lower-energy magnons within a d-wave-shaped “island of stability” around the Brillouin zone origin, g-wave and i-wave splittings cannot. We identify quasi-two-dimensional d-wave altermagnets with negligible spin-orbit interaction as ideal platforms for exploring altermagnetic quantum spin dynamics, and predict experimentally resolvable magnon damping in organic altermagnets. These findings establish a quasiparticle stability trichotomy of nonrelativistic collinear magnets, with the three principal cases shown in Fig. 1, highlighting how altermagnets depart from established paradigms in quantum magnetism and spontaneous quasiparticle decay.
Trichotomy of magnon stability in collinear magnets — kinematics
Nonrelativistic collinear magnets can be categorized into three distinct classes using spin space group arguments, which involve the partial-wave decomposition of the spin splitting in their electronic or magnonic bands [19, 20, 22], see Fig. 1. (i) s-wave spin splitting: Magnets in this class, such as ferromagnets and ferrimagnets, exhibit a finite net magnetic moment in real space. (ii) No spin splitting: Characterized by compensated order in real space, these magnets exhibit oppositely ordered sublattices connected by inversion or translation symmetry and are called antiferromagnets. (iii) Even-parity spin splitting beyond s-wave (d-, g-, or i-wave): These magnets—now recognized as altermagnets [19]—display compensated order in real space, but their oppositely ordered sublattices are not related by inversion or translation symmetry but instead, e.g., by rotational symmetry.
For insulators with localized electronic degrees of freedom, all three classes of magnets adhere to the nonrelativistic SO$(3)$-symmetric Heisenberg model ($\hbar=1$),
$H=\frac{1}{2}\sum_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}J_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}{\boldsymbol{S}}_{{\boldsymbol{r}}}\cdot{\boldsymbol{S}}_{{\boldsymbol{r}}^{\prime}},$ (1)
where ${\boldsymbol{S}}_{\boldsymbol{r}}$ is a local spin operator of length $S$ at site ${\boldsymbol{r}}$, and $J_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}$ the exchange constant between spins at ${\boldsymbol{r}}$ and ${\boldsymbol{r}}^{\prime}$. In the collinear ground state, spontaneous symmetry breaking reduces SO$(3)$ to SO$(2)$, with the ordered moment chosen along the $z$-axis. As a result, Goldstone modes emerge, and the total spin component $S^{z}=\sum_{\boldsymbol{r}}S_{\boldsymbol{r}}^{z}$ remains a good quantum number. Single spin-flip excitations (magnons) thus carry spin ±$\pm 1$ (also denoted by $\uparrow$ and $\downarrow$) relative to the ground state, and follow dispersions $\epsilon_{\pm}({\boldsymbol{k}})$, where ${\boldsymbol{k}}$ is the crystal momentum, and “$\pm$” indicates spin, not necessarily energetic order. Optical magnon branches are neglected as they do not impact low-energy physics.
Although spin is conserved, magnon number is not—except in ferromagnets—allowing for quantum many-body interactions that change magnon count while preserving total spin [8, 3]. Among these, three-magnon processes are forbidden by spin conservation, making four-magnon interactions the lowest-order relevant scattering mechanisms [8]. These fall into three categories: (1) Two-in-two-out, which require thermally excited magnons and are inactive at zero temperature; (2) None-in-four-out (and vice versa), which violate energy conservation and thus do not affect stability; (3) One-in-three-out (and vice versa), the only processes relevant at zero temperature. For instance, an initial spin-$\uparrow$ magnon can decay into two spin-$\uparrow$ magnons and one spin-$\downarrow$ magnon, conserving total spin. The feasibility of such a three-magnon decay depends on energy and momentum conservation:
$\epsilon_{\pm}({\boldsymbol{k}})=\epsilon_{\pm}({\boldsymbol{q}})+\epsilon_{\pm}({\boldsymbol{p}})+\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{q}}-{\boldsymbol{p}}).$ (2)
Thus, magnon stability in collinear nonrelativistic magnets at absolute zero reduces to determining whether Eq. (2) has solutions. If it does, decays occur, provided the scattering matrix element is nonzero.
Class-(i) magnets: Ferromagnets inherently support stable magnons because their ground state is an exact eigenstate. Ferrimagnets, however, fluctuate because their sublattice-polarized ground state is not an eigenstate. Having two antiparallel but uncompensated sublattice magnetizations, they feature two magnon branches: a gapless branch, $\epsilon_{-}({\boldsymbol{k}})=A|{\boldsymbol{k}}|^{2}$, and a gapped branch, $\epsilon_{+}({\boldsymbol{k}})=A|{\boldsymbol{k}}|^{2}+\Delta_{0}$, where $A>0$ is the spin-wave stiffness and $\Delta_{0}>0$ the gap, see Fig. 1a. Here, s-wave spin splitting manifests as an isotropic energy difference: $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=\Delta_{0}$. The positive curvature of the gapless branch implies that any initial magnon could theoretically decay into three magnons at smaller momenta while conserving energy and momentum. However, Eq. (2) requires at least one decay product to have opposite spin, precluding decay into the gapless branch alone. Decay becomes possible only at energies exceeding $\Delta_{0}$, where one decay product can transition into the gapped branch. Consequently, the gapless magnon branch in class-(i) magnets is universally stable below the energy gap $\Delta_{0}$ and protected by s-wave spin splitting. The gapped branch, however, can always decay.
Class-(ii) magnets: In antiferromagnets, which lack spin splitting, the magnon branches are degenerate: $\epsilon_{+}({\boldsymbol{k}})=\epsilon_{-}({\boldsymbol{k}})=\epsilon({\boldsymbol{k}})$. Magnon stability at long wavelengths depends on the curvature of the dispersion relation. Consider the dispersion $\epsilon({\boldsymbol{k}})=v|{\boldsymbol{k}}|+\kappa|{\boldsymbol{k}}|^{3}$ where $v>0$ is the magnon velocity and $\kappa$ characterizes the leading nonlinearity. The sign of $\kappa$ determines the stability [3]: For $\kappa>0$, the magnon dispersion has positive curvature, in which case Eq. (2) has solutions and magnons are unstable. If $\kappa<0$, Eq. (2) has no solutions and magnons are stable. Beyond the isotropic approximation, one has to account for lattice symmetries, that is, for $\kappa$ exhibiting directional dependence, resulting in magnon damping varying with orientation. Thus, class-(ii) magnets can support stable gapless magnons under specific conditions dictated by leading nonlinearities of the dispersion, but lack spin splitting to universally protect magnons from decay.
Figure 2: Kinematics of magnon decays in class-(iii) magnets: altermagnets. Low-energy dispersion of spin-$\uparrow$ and spin-$\downarrow$ magnons with energy $\epsilon_{+}({\boldsymbol{k}})$ and $\epsilon_{-}({\boldsymbol{k}})$, respectively. Along a nodal plane, the magnon dispersion is degenerate. (a) Decay channel of nodal-plane magnons. They decay into a magnon close to the origin and two magnons of opposite spin in the lower branch to the left and right of the nodal line, respectively. (b) Decay channel of the lower-energy magnon. It decays into a magnon close to the origin and two magnons with the same spin to its left and right, respectively. (c) First decay channel of the higher-energy magnon. It decays into a magnon close to the origin and two magnons with the same spin to its left and right, respectively. (d) Second decay channel of the higher-energy magnon. It decays into two magnons of the same spin close to the origin and a magnon in the lower-energy branch.
Class-(iii) magnets: To derive general insights into magnon decay in altermagnets, we perform a low-energy expansion of the magnon dispersion in two dimensions. For simplicity, we assume the dispersion to be spatially isotropic, allowing crystallographic symmetries to enter only through the altermagnetic splitting. Under these approximations, the magnon dispersion takes the form
$\epsilon_{\pm}({\boldsymbol{k}})=v|{\boldsymbol{k}}|+\kappa|{\boldsymbol{k}}|^{3}\pm\Delta\sin\left(n\varphi_{{\boldsymbol{k}}}\right)|{\boldsymbol{k}}|^{n},$ (3)
where $v>0$ is the magnon velocity, $\kappa$ represents the leading (cubic) non-altermagnetic nonlinearity, $\tan\varphi_{{\boldsymbol{k}}}=k_{y}/k_{x}$, and $\Delta\geq 0$ parametrizes the altermagnetic splitting with a d-wave ($n=2$), g-wave ($n=4$) or i-wave character ($n=6$). The unconventional time-reversal symmetry breaking results in an even-parity spin splitting gap $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=2\Delta\sin\left(n\varphi_{{\boldsymbol{k}}}\right)|{\boldsymbol{k}}|^{n}$ that modulates with $n\varphi_{{\boldsymbol{k}}}$. Along a nodal plane (or line in 2D), e.g., ${\boldsymbol{k}}=(k_{x},k_{y})=(k,0)$, the splitting vanishes, and the degenerate spectrum suggests that the curvature argument relevant to class-(ii) magnets will also play a role here. Away from the nodal planes, the magnon bands split, opening up novel decay channels, see Fig. 1c.
We begin with discussing magnon stability along the nodal planes. If $\kappa>0$, nodal-plane magnons can decay similarly to magnons in class-(ii) magnets. This leaves $\kappa<0$ as the relevant case, where magnons in class-(ii) magnets remain stable. Consider the case depicted in Fig. 2a, where a nodal plane magnon decays into a magnon close to the Goldstone mode, and two lower-branch magnons to the left and right of the nodal plane. As shown in the Supplementary Note 1, this process obeys Eq. (2), if $\Delta>\Delta^{\star}$, where
$\Delta^{\star}=\sqrt{3}\frac{2^{n-5/2}}{nk^{n-2}}\sqrt{-\kappa}\sqrt{4v+3k^{2}\kappa}\quad\text{for}\quad\kappa<0.$ (4)
In the long-wavelength limit ($k\to 0$), $\Delta^{\star}$ diverges for $n=4$ (g-wave) and $n=6$ (i-wave) due to the $k^{n-2}$ factor in the denominator. This indicates that nodal-plane magnons are stable in the vicinity of the origin. For $n=2$ (d-wave), however, $\Delta^{\star}$ converges to
$\Delta^{\star}_{\text{d-wave}}=\sqrt{-\frac{3}{2}v\kappa}\quad\text{for}\quad\kappa<0.$ (5)
Thus, d-wave magnets with a negatively curved magnon spectrum ($\kappa<0$) may still exhibit unstable nodal-plane magnons if the altermagnetic splitting exceeds $\Delta^{\star}_{\text{d-wave}}$.
Next, we analyze the decay of the lower-branch magnon along a general direction, where it could decay into two lower-branch magnons and a magnon with opposite spin close to the Goldstone mode, as shown in Fig. 2b. For $\kappa<0$, this decay process is generally kinematically forbidden. For $\kappa>0$, as shown in Supplementary Note 1, it remains forbidden if
$|{\boldsymbol{k}}|^{3-n}<\frac{2^{2-n}\left(2^{n}-2\right)}{3\kappa}\Delta\left|\sin(n\varphi_{{\boldsymbol{k}}})\right|.$ (6)
Thus, in d-wave altermagnets ($n=2$), lower-branch magnons at $|{\boldsymbol{k}}|<k^{\star}_{\text{d-wave}}\left|\sin\left(2\varphi_{{\boldsymbol{k}}}\right)\right|$, with
$k^{\star}_{\text{d-wave}}=\frac{2\Delta}{3\kappa}\quad\text{for}\quad\kappa>0$ (7)
are stable. No such long-wavelength solution can be found for g-wave magnets ($n=4$) and i-wave magnets ($n=6$) because of the negative power of $k^{3-n}$, which turns the inequality into a statement about short wavelengths, where the approximation in Eq. (3) is no longer valid. Thus, the altermagnetic splitting of g-wave and i-wave magnets does not protect the lower branch from decays and it is the dominating cubic nonlinearity that decides the fate of their lower-branch magnons.
These qualitative kinematic considerations for the lower branch can be summarized as follows. Two cases must be distinguished for d-wave magnets: For $\kappa<0$ and $\Delta<\Delta^{\star}_{\text{d-wave}}$, the lower magnon branch remains stable. For $\kappa<0$ and $\Delta>\Delta^{\star}_{\text{d-wave}}$ or for $\kappa>0$, it exhibits a d-wave-shaped “island of stability” around the origin, with nodes at the nodal planes. The extent of this stability region is given by $k^{\star}_{\text{d-wave}}$ in Eq. (7). In contrast, in g-wave and i-wave altermagnets, its stability depends solely on the sign of $\kappa$, as in class-(ii) magnets. The qualitative distinction between d-wave and g,i-wave altermagnets arises because the altermagnetic splitting in d-wave magnets depends on a lower power of momentum than the leading (cubic) nonaltermagnetic nonlinearity, $\kappa$. In contrast, for g-wave and i-wave altermagnets, the altermagnetic splitting becomes irrelevant in the long-wavelength limit compared to the cubic nonlinearities  [^1] In fine-tuned systems with $\kappa=0$, which we do not explore here, quintic terms, $\sim|{\boldsymbol{k}}|^{5}$, become the dominant nonaltermagnetic nonlinearity, and the quartic g-wave altermagnetic splitting becomes relevant in the long-wavelength limit. Similarly, if both cubic and quintic terms vanish, septic nonlinearities, $\sim|{\boldsymbol{k}}|^{7}$, get dominated by the sextic i-wave altermagnetic splitting..
We now turn to the upper magnon branch, for which two decay channels are indicated in Fig. 2c and Fig. 2d. In the process in Fig. 2c, an upper-branch magnon decays into a mode close to the origin and two modes to the left and right. The process is possible in d-wave magnets even for $\kappa<0$ because their altermagnetic splitting dominates over the cubic nonlinearity at long wavelengths. In contrast, g-wave and i-wave altermagnets do not support this process. However, the process in Fig. 2d has the upper-branch magnon decay into a lower-branch magnon at approximately the same momentum and two upper-branch magnons around the origin. As this decay is always kinematically possible, the upper magnon branch in altermagnets is universally unstable.
Nonlinear spin-wave theory of altermagnets
Figure 3: Nonlinear spin-wave analysis of a two-dimensional d-wave altermagnet. (a) Schematic of the spin model for a two-dimensional d-wave altermagnet on a checkerboard lattice, described by the Hamiltonian in Eq. (8). The system hosts a collinear $(\pi,\pi)$ Néel order with nearest-neighbor (white bonds) and next-nearest-neighbor (red and blue bonds) exchange interactions $J_{1}$ and $J_{2}\pm\Delta$, respectively. The checkerboard modulation—parametrized by $\Delta$—arises from nonmagnetic atoms (silver spheres) that alter the diagonal exchange, inducing a $d$-wave spin splitting of magnons. (b) Diagrammatic representation of the three-magnon scattering process, the leading mechanism for magnon decay. Due to momentum conservation (up to a reciprocal lattice vector), the sum of the momenta of the three intermediate states equals the initial momentum ${\boldsymbol{k}}$. Total spin conservation ensures that an initial single spin-$\uparrow$ magnon scatters into two spin-$\uparrow$ magnons and one spin-$\downarrow$ magnon, with a similar process for an initial spin-$\downarrow$ magnon. (c-f) Renormalized magnon spectrum (lines with dot markers) compared to the linear magnon spectrum (solid lines) for $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$, and selected values of $\Delta$ as indicated. The altermagnetic splitting is most pronounced at $(\frac{\pi}{2},\pm\frac{\pi}{2})$ and zero along the nodal planes, e.g., in the direction from $(0,0)$ to $(\pi,0)$. (g-j) Magnon damping $\Gamma_{\pm}({{\boldsymbol{k}}})$, highlighting the instability of the upper magnon branch. The inset in panel (g) shows the magnetic Brillouin zone with indicated high-symmetry paths. The color of these paths in used on the horizontal momentum axes in panels (c-j) to indicate direction. The cyan/magenta checkerboard pattern indicates the sign of the d-wave altermagnetic spin splitting. The legend in panel (j) distinguishes between nonlinear spin-wave (NLSW) and linear spin-wave (LSW) calculations.
Figure 4: Magnon decay phase space analysis and long-wavelength scaling of magnon decay rate in d-wave altermagnets. (a) Three-magnon density of states $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$, encoding the available decay phase space for the spin-$\uparrow$ magnon with energy $\epsilon_{+}({\boldsymbol{k}})$ (magenta line) for moderate altermagnetic splitting ($\Delta=0.3$). (b) Same as (a) but for $\Delta=1.0$, pushing the upper magnon branch deeper into the continuum. The colors on the horizontal momentum axes in (a) and (b) are identical to those in the Brillouin zone inset in Fig. 3g. Parameters used in (a) and (b) are $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$ and $\Delta$ as indicated, realizing the case of a negative cubic nonlinearity: $\kappa<0$. (c) On-shell three-magnon density of states for the lower-energy magnon branch, $\mathcal{D}_{\xi}(\epsilon_{\xi}({\boldsymbol{k}}),{\boldsymbol{k}})$, in vicinity of the Brillouin zone origin. $\xi$ is chosen so that the lower band is picked. The d-wave-shaped “island of stability” agrees with the indicated $k^{\star}_{\text{d-wave}}$ as obtained from the kinematic analysis in Eq. (7). The nodal plane is indicated as a dotted diagonal line. Parameters are $J_{1}=1/10$, $J_{2}=-1$, $S=1/2$, and $\Delta=0.5$, realizing the case of a positive cubic nonlinearity: $\kappa>0$. (d) Long-wavelength scaling of the magnon decay rate $\Gamma$ along selected directions. The upper magnon branch shows a characteristic $|{\boldsymbol{k}}|^{5}$ scaling away from the nodal planes that dominates potential subleading $|{\boldsymbol{k}}|^{7}$ contributions. Along the nodal planes, the $|{\boldsymbol{k}}|^{7}$ scaling is the leading contribution. The indicated scaling assumes two-dimensional d-wave altermagnets.
To expand upon the preceding qualitative discussion of magnon decay we consider a d-wave altermagnet on the checkerboard lattice (see Fig. 3a), for which we carry out a nonlinear spin-wave analysis. The spin Hamiltonian under investigation is given by
$H=\sum_{\expectationvalue{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}}J_{1}{\boldsymbol{S}}_{{\boldsymbol{r}}}\cdot{\boldsymbol{S}}_{{\boldsymbol{r}}^{\prime}}+\sum_{\expectationvalue{\expectationvalue{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}}_{\pm}}(J_{2}\pm\Delta){\boldsymbol{S}}_{\boldsymbol{r}}\cdot{\boldsymbol{S}}_{{\boldsymbol{r}}^{\prime}}.$ (8)
Here, $J_{1}>0$ is the antiferromagnetic nearest-neighbor, and $J_{2}$ the second-nearest-neighbor exchange interaction. $\Delta>0$ parametrizes the d-wave altermagnetic magnon band splitting. Its sign alternates between white and black squares of the checkerboard, as indicated in Fig. 3a.
Using a $1/S$ spin-wave expansion around the Néel-ordered ground state based on the Holstein-Primakoff transformation [25], we derive the noninteracting single-magnon spectrum $\epsilon_{\pm}({\boldsymbol{k}})=S\sqrt{A({\boldsymbol{k}})^{2}-B({\boldsymbol{k}})^{2}}\pm S\Delta({\boldsymbol{k}})$, with $A({\boldsymbol{k}})$, $B({\boldsymbol{k}})$, and $\Delta({\boldsymbol{k}})$ given in Methods. At low energies, the spectrum can be mapped onto Eq. (3) up to a rotation of the nodal planes and an additional directional dependence of the cubic nonlinearits. Up to order $1/S^{2}$, the renormalized single-magnon spectrum is given by $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})=\epsilon_{\pm}({\boldsymbol{k}})+\delta\epsilon_{\pm}({\boldsymbol{k}})-\mathrm{i}\Gamma_{\pm}({\boldsymbol{k}})$. The real $\delta\epsilon_{\pm}({\boldsymbol{k}})$ corrects the magnon energy and the magnon lifetime $\tau_{\pm}({\boldsymbol{k}})=-1/\mathrm{Im[\tilde{\epsilon}_{\pm}({\boldsymbol{k}})]}$ derives from the decay rate
$\Gamma_{\pm}({\boldsymbol{k}})=$ $\frac{\pi}{N_{\textrm{muc}}^{2}}\sum_{{\boldsymbol{p}},{\boldsymbol{q}}}|W_{\pm}({\boldsymbol{k}},{\boldsymbol{p}},{\boldsymbol{q}})|^{2}$
×$\times\delta\left(\epsilon_{\pm}({\boldsymbol{k}})-\epsilon_{\pm}({\boldsymbol{p}})-\epsilon_{\pm}({\boldsymbol{q}})-\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}})\right),$ (9)
associated with the imaginary part of the self-energy diagram in Fig. 3b, where one encounters the energy conservation condition of Eq. (2). In Eq. (9), $N_{\textrm{muc}}$ is the number of magnetic unit cells, the momenta sums run over the magnetic Brillouin zone, and $W_{\pm}({\boldsymbol{k}},{\boldsymbol{p}},{\boldsymbol{q}})$ is the one-in-three-out interaction vertex (see Methods and Supplementary Note 2).
The interacting magnon spectrum $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})$ is shown in Figs. 3c-j for parameters realizing $\kappa<0$. It exhibits clear d-wave splitting, which increases as $\Delta$ increases from Fig. 3c to Fig. 3f. The many-body interactions cause a stiffening of both bands and, importantly, a damping of the upper magnon branch, see Figs. 3g-j. The damping grows with increasing $\Delta$ and is most pronounced in the short-wavelength limit. Since $\Delta>\Delta^{\star}_{\text{d-wave}}$ in Figs. 3h, the magnon damping is also finite along the nodal planes, that is, along the direction from $(0,0)$ to $(\pi,0)$, in agreement with the kinematic predictions.
The main features of $\Gamma_{\pm}({\boldsymbol{k}})$ in Figs. 3g-j can be understood by setting the interaction vertex in Eq. (9) to one, suggesting that $\Gamma_{\pm}({\boldsymbol{k}})\propto\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})$, where
$\mathcal{D}_{\pm}(\omega,{\boldsymbol{k}})=\frac{1}{N_{\textrm{muc}}^{2}}\sum_{{\boldsymbol{p}},{\boldsymbol{q}}}\delta\left(\omega-\epsilon_{\pm}({\boldsymbol{p}})-\epsilon_{\pm}({\boldsymbol{q}})-\epsilon_{\mp}({\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}})\right)$ (10)
is the density of states of the three-magnon continuum built from three magnons whose spin adds up to ±$\pm 1$. Carrying a finite net spin, the $\mathcal{D}_{\pm}(\omega,{\boldsymbol{k}})$ continua also exhibit the d-wave splitting. Whenever these continua overlap with the single-magnon branch of the same spin, that is, whenever $\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})\neq 0$, spontaneous decays are kinematically possible [3].
Figure 4a shows $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$ together with the noninteracting magnon spectrum for $\kappa<0$ and $\Delta=0.3$. The spin-$\uparrow$ magnon with energy $\epsilon_{+}({\boldsymbol{k}})$ is embedded within the $\mathcal{D}_{+}(\omega,{\boldsymbol{k}})$ continuum along the direction $(0,0)$ to $(\frac{\pi}{2},-\frac{\pi}{2})$, and can therefore decay. In contrast, along the direction $(0,0)$ to $(\frac{\pi}{2},\frac{\pi}{2})$ the continuum begins exactly at that branch’s energy (due to the Goldstone mode), precluding decays. A similar discussion applies to the spin-$\downarrow$ magnon with energy $\epsilon_{-}({\boldsymbol{k}})$ and the $\mathcal{D}_{-}(\omega,{\boldsymbol{k}})$ continuum (not shown). Along the nodal-plane direction from $(0,0)$ to $(\pi,0)$, the continuum also begins exactly at the single-particle energies because $\Delta<\Delta^{\star}_{\text{d-wave}}$, rendering magnons stable.
For $\kappa<0$ and $\Delta=1.0$, the lower magnon branch exhibits a local minimum at $(-\frac{\pi}{2},\frac{\pi}{2})$, which causes a rapid increase in the three-particle DOS at about three times its energy, as shown in Fig. 4b. As a result, $\Gamma({\boldsymbol{k}})$ of the upper branch also exhibits a sharp increase at this energy, in agreement with what is seen in Fig. 3j. Since $\Delta>\Delta^{\star}_{\text{d-wave}}$, the magnon energies are embedded within the continua along the nodal-plane direction from $(0,0)$ to $(\pi,0)$, explaining the finite decay rate seen in Fig. 3j.
For $\kappa>0$, kinematics predicts an island of stability of size $k^{\star}_{\text{d-wave}}$ for the lower magnon branch around the origin. Indeed, as shown in Fig. 4c, the relevant continua for the lower magnon mode exhibit a pronounced drop to zero with d-wave-shaped lobes. Around the origin, the density of states of the continuum stays finite only along the nodal planes, where the altermagnetic splitting cannot protect the lower branch from decays.
Long-wavelength scaling of magnon decay rates in altermagnets
The existence of Goldstone modes implies universal low-energy physics, raising the question how $\Gamma_{\pm}({\boldsymbol{k}})$ in Eq. (9) scales with $|{\boldsymbol{k}}|$ as $|{\boldsymbol{k}}|\to 0$. Its scaling depends on both that of the available decay phase space, i.e., the scaling of the on-shell three-magnon density of states $\mathcal{D}_{\pm}(\epsilon_{\pm}({\boldsymbol{k}}),{\boldsymbol{k}})$—as given by Eq. (10) with $\omega\to\epsilon_{\pm}({\boldsymbol{k}})$—and that of $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})$.
We begin with the decay phase space, which we call $V({\boldsymbol{k}})$ for brevity. At long wavelengths, the linear magnon spectrum $\epsilon({\boldsymbol{k}})=v|{\boldsymbol{k}}|$ is the reference point. A magnon at ${\boldsymbol{k}}$ is “ready to decay” into magnons at ${\boldsymbol{p}}$, ${\boldsymbol{q}}$, and ${\boldsymbol{k}}-{\boldsymbol{p}}-{\boldsymbol{q}}$ along the line from the origin to ${\boldsymbol{k}}$. Hence, any positive curvature—be it due to a quadratic or cubic nonlinearity—activates decays of the type in Fig. 2c in a threshold behavior. Since the two momenta ${\boldsymbol{q}}$ and ${\boldsymbol{p}}$ have to be almost aligned with ${\boldsymbol{k}}$ to fulfill energy and momentum conservation, the four-dimensional decay phase space has a cigar shape. In the two long directions, it is constrained by $|{\boldsymbol{k}}|$ and in the other two short directions by $|{\boldsymbol{k}}|^{3/2}$. It follows that $V({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ for a positively curved magnon dispersion. In case of altermagnets with their even-parity spin split dispersion $\epsilon_{\pm}({\boldsymbol{k}})=v|{\boldsymbol{k}}|\pm\Delta\sin(n\varphi_{{\boldsymbol{k}}})|{\boldsymbol{k}}|^{n}$, this observation implies that $V({\boldsymbol{k}})$ of the upper branch is zero at $\Delta=0$ but jumps to a finite value at $\Delta\neq 0$.
The threshold value of $\Delta$ becomes nonzero and $n$-dependent once additional nonlinearities are taken into account. Including a negative cubic nonlinearity, $\kappa<0$, as in Eq. (3), we find that $\Delta$ has to exceed $-C_{n}\kappa|{\boldsymbol{k}}|^{3-n}$, with $C_{n}>0$ being a constant, to cause decays of the upper branch at $|{\boldsymbol{k}}|$. For d-wave magnets ($n=2$), we find that $\Delta>-C_{n}\kappa|{\boldsymbol{k}}|$ holds for any value of $\kappa<0$ if only $|{\boldsymbol{k}}|$ is small enough. Thus, upper-branch magnons in d-wave magnets have a decay phase space $V_{\text{d-wave}}({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ for $\kappa<0$ as $|{\boldsymbol{k}}|\to 0$. In contrast, g-wave and i-wave magnets have a negative power of $|{\boldsymbol{k}}|^{3-n}$, implying that their threshold value for $\Delta$ diverges as $|{\boldsymbol{k}}|\to 0$. Thus, the upper magnon branch of g-wave and i-wave magnets with $\kappa<0$ does not exhibit the $V({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{3}$ long-wavelength scaling. Instead, other higher-order-in-momentum mechanisms have to be identified; an example is the mechanism depicted in Fig. 2d.
With the relevant decay process for d-wave altermagnets identified, we plug the “cigar-shaped” scaling ($\propto|{\boldsymbol{k}}|$ for the two long and $\propto|\Delta|^{1/2}|{\boldsymbol{k}}|^{3/2}$ for the two short directions) for the respective components of ${\boldsymbol{p}}$ and ${\boldsymbol{q}}$ into $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})$ and take the limit $|{\boldsymbol{k}}|\to 0$. We find $W_{\pm}({\boldsymbol{k}},{\boldsymbol{q}},{\boldsymbol{p}})\propto|\Delta||{\boldsymbol{k}}||\sin(2\varphi_{\boldsymbol{k}})|$, where we have restored the angular dependence. Therefore, by combining the square of the vertex with the decay phase space scaling according to Eq. (9), the damping of the upper magnon branch in d-wave altermagnets is given by
$\Gamma_{\text{upper}}({\boldsymbol{k}})\propto|\Delta|^{2}|{\boldsymbol{k}}|^{5}|\sin(2\varphi_{\boldsymbol{k}})|^{2}\quad\text{as}\quad|{\boldsymbol{k}}|\to 0,$ (11)
as indicated in Fig. 4d.
Magnons along the nodal planes are stable if $\Delta<\Delta^{\star}_{\text{d-wave}}$ but exhibit the decays shown in Fig. 2a for $\Delta>\Delta^{\star}_{\text{d-wave}}$, with $\Delta^{\star}_{\text{d-wave}}$ given in Eq. (5). In this case, the cigar-shaped decay space has a short-axes scaling of $|{\boldsymbol{k}}|^{2}$, such that the damping along the nodes is given by
$\Gamma_{\text{node}}({\boldsymbol{k}})\propto\Theta\left(\Delta-\Delta^{\star}_{\text{d-wave}}\right)|{\boldsymbol{k}}|^{7}\quad\text{as}\quad|{\boldsymbol{k}}|\to 0,$ (12)
which is also indicated in Fig. 4d. $\Theta$ indicates the Heaviside step function. Also, for $\kappa>0$, an isotropic scaling $\Gamma({\boldsymbol{k}})\propto|{\boldsymbol{k}}|^{7}$ is expected [11].
We conclude that in the case of d-wave altermagnets, the cubic nonlinearity is subleading in the long-wavelength limit and away from the nodal planes, giving rise to the unique leading $|{\boldsymbol{k}}|^{5}$ scaling of the decay rate in Eq. (11). For further details on the long-wavelength scaling of magnon damping, see Supplementary Notes 3 and 4. We emphasize that our results hold for two-dimensional altermagnets.
Nonperturbative quantum simulations of altermagnets
Figure 5: Comparison of nonperturbative and perturbative transverse quantum dynamics in two-dimensional d-wave altermagnets. (a) The transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$, as obtained from DMRG$+$tMPO, shows clear altermagnetic splitting along the path $(0,0)$ to $(\frac{\pi}{2},\frac{\pi}{2})$ in reciprocal space. Around this point, the spectral weight of the upper magnon branch gets almost completely wiped out due to hybridization with continua. A small XXZ anisotropy $\lambda=0.02$ was added for numerical stability. It causes a tiny gap of the Goldstone modes. Note the double linear color bar scale. Dashed white lines indicate points at which simulation results for different simulation geometries were patched together (see Methods). As the simulation on cylinders weakly breaks the exact four-fold rotational symmetry of the lattice, small discontinuities in the intensity can be identified. (b) The single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)$ as obtained within perturbative nonlinear spin-wave (NLSW) theory of $1/S^{2}$ accuracy modulo additional off-shell corrections. The colors on the horizontal momentum axes in (a) and (b) match the colors of the high-symmetry paths indicated in the Brillouin zone in the inset of (b). (c,d) Line cuts of the nonperturbative $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ versus frequency at momenta $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$, respectively. (e,f) Line cuts of the perturbative $\mathcal{A}({\boldsymbol{k}},\omega)$ versus frequency at momenta $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$, respectively. Parameters read $J_{1}=1$, $J_{2}=-0.5$, $S=1/2$, and $\Delta$ as indicated.
The observation of magnon decays arising from the nonrelativistic altermagnetic beyond-s-wave band splitting was derived within the framework of perturbation theory. At low energies, where both the decay phase space and magnon-magnon interaction vertices are small, perturbation theory provides reliable results [8, 9]. At higher energies, our calculations suggest $\Gamma/\epsilon\lesssim 0.1$ for $S=1/2$, indicating that magnons remain well-defined, albeit damped, quasiparticles, validating the perturbative approach. However, small spin $S$, non-universal high-energy decays, and enhanced quantum fluctuations in low-dimensional systems necessitate nonperturbative methods to fully assess magnon stability. We address this using density matrix renormalization group (DMRG) [26, 27] and matrix product operator time evolution (tMPO) [28] techniques to extract the dynamical spin structure factor $\mathcal{S}_{\mu\nu}({\boldsymbol{k}},\omega)$ (with $\mu$ and $\nu$ denoting $x$, $y$, and $z$), as measured in inelastic neutron scattering (see Methods for details).
Figure 5a presents the transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)=\mathcal{S}_{xx}({\boldsymbol{k}},\omega)+\mathcal{S}_{yy}({\boldsymbol{k}},\omega)$ from nonperturbative simulations, revealing distinct magnon peaks. Peak intensity decreases due to magnon damping and non-Bravais lattice form factors, explaining the vanishing intensity at $(0,0)$. The lower mode remains sharp (within Gaussian filtering of width $\sigma_{\omega}=0.0896$), while the upper mode broadens along $(0,0)\to(\frac{\pi}{2},\frac{\pi}{2})$, aligning with perturbation theory. At $(\frac{\pi}{2},\frac{\pi}{2})$, its non-Lorentzian shape at large $\Delta$ signals strong nonperturbative effects.
We find consistently good agreement between the nonperturbative spectrum in Fig. 5a and the $1/S^{2}$ off-shell single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)$ in Fig. 5b (see Methods). Linecuts through $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ and $\mathcal{A}({\boldsymbol{k}},\omega)$ in Fig. 5c,d and Fig. 5e,f at $(\frac{\pi}{2},\frac{\pi}{2})$ and $(\pi,0)$ show remarkable agreement. However, perturbation theory underestimates nonperturbative high-energy tails (compare Fig. 5d with Fig. 5f), possibly due to three-particle contributions, spinon continua [29], or Higgs resonances [30]. Capturing the full quantum dynamics of $S=1/2$ altermagnets may thus require incorporating amplitude oscillations of the order parameter, which lie beyond perturbation theory.
Experimental and theoretical implications
The isotropic Heisenberg model, with its spin-rotation symmetry, provides a reasonable approximation for quantum magnets when relativistic effects are negligible. In real materials, relativistic corrections such as magnetocrystalline anisotropy, $\lambda\sum_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}S_{{\boldsymbol{r}}}^{z}S_{{\boldsymbol{r}}^{\prime}}^{z}$, introduce perturbations that affect magnon stability. For easy-axis anisotropy ($\lambda<0$), relevant for altermagnetic candidates like MnF2, CoF2, and $\alpha$-Fe2O3 [19, 22], both Goldstone modes acquire a gap, growing as $\sqrt{\lambda J}$ due to the exchange enhancement. As $\lambda$ increases, the decay phase space shrinks, and beyond a critical threshold $\lambda^{\star}$, all magnons become stable, as expected in the Ising limit where the Néel ground state is an exact eigenstate. This provides a straightforward experimental criterion: for instance, neutron scattering data [31] set an upper bound of $\sim 120\,\mu$eV for the magnon splitting of MnF2, while its long-wavelength gap exceeds 111\,1meV, precluding three-magnon decays.
[…]
Beyond global anisotropy, relativistic interactions such as Dzyaloshinskii-Moriya and pseudo-dipolar exchange can break spin-rotation symmetry, destabilizing magnetic order and enabling additional two-magnon decay processes [12, 13, 14, 15, 16, 17]. External magnetic fields can induce similar effects [10, 18]. These interactions contribute to magnon damping already at order $1/S$ [3], potentially overshadowing intrinsic $1/S^{2}$ many-body effects characteristic of altermagnets.
For an ideal platform to study pristine altermagnetic many-body magnon physics, the material should exhibit minimal relativistic corrections, or at most weak easy-axis anisotropy, to preserve approximate spin conservation. A quasi-two-dimensional structure and small spin quantum number $S$ is desirable to enhance quantum fluctuations, and a d-wave altermagnetic character leads to the most pronounced effects. The material should also be insulating to suppress electron-magnon damping, which also exhibits even-parity symmetries beyond s-wave [22, 33].
[…]
In discussions of altermagnets, their dual nature is often emphasized: they combine features of ferromagnets—such as spin splitting and time-reversal symmetry breaking—with those of antiferromagnets, including a compensated net moment [47]. The stability of the gapless, long-wavelength magnons discussed here reflects this hybrid character. Class-(i) magnets stabilize their gapless magnons via s-wave spin splitting, but their gapped branch is universally unstable. In contrast, class-(ii) magnets lack protective spin splitting, yet the curvature of their magnon dispersion can ensure stability under specific conditions. Class-(iii) magnets—altermagnets—present a combination of these mechanisms. At nodal planes in momentum space, where spin splitting vanishes, their magnons behave like those of class-(ii) magnets, albeit with corrections for d-wave altermagnets. However, in directions with spin-split magnon bands, the lower branch is stabilized in a manner reminiscent of class-(i) magnets for d-wave altermagnets, but not for g-wave and i-wave cases. As in class-(i) magnets, the upper branch remains universally unstable for that direction.
Beyond their significance in magnetism, altermagnets occupy a distinctive position in the broader landscape of spontaneous quasiparticle decay. Their continuous SO$(2)$ spin-space rotation symmetry makes spin-conserving three-particle decay the dominant instability, while two-particle decay is typically the main mechanism in other systems, such as phonons [5, 7] or magnons in relativistic and noncollinear magnets [3]. In collinear magnets, two-particle decay can also be suppressed by a discrete $C_{2}$ spin rotation symmetry [11, 3, 48], but in this case, three-particle decays are not constrained to conserve spin. Intriguingly, even continuous spin-space symmetries do not always suppress two-particle decay in all types of magnetic quasiparticles. For example, in quantum spin-gap magnets with singlet ground states, two-triplon decay remains allowed despite spin-rotation symmetry, as antisymmetric two-triplon combinations can still conserve total spin [49]. In these systems, additional real-space symmetries, such as reflections, may be necessary to suppress two-particle decays [50, 3]. Therefore, altermagnets present an intriguing platform for studying symmetry-constrained many-body dynamics.
Furthermore, while systems with multiple Goldstone modes—such as phonons in crystals or magnons in noncollinear magnets [3]—typically allow decay from faster to slower modes, this is generally driven by differences in their linear dispersion, for instance, by the velocity mismatch between transverse and longitudinal phonons. In altermagnets, however, the splitting between the modes is at least quadratic (d-wave), fundamentally altering the long-wavelength decay kinematics. Additionally, the even-parity nature of this splitting introduces a unique “island of stability” in magnon lifetimes. These findings highlight the exceptional many-body physics of altermagnets, signaling a marked departure from conventional paradigms in quantum magnetism and quasiparticle decay in condensed matter systems.
Note added
While preparing the manuscript we have become aware of independent complementary work on spontaneous magnon decay in d-wave altermagnets by Cichutek et al. [51], where an analytical long-wavelength analysis is presented for the checkerboard model.
Acknowledgements.
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project No. 504261060 (Emmy Noether Programme) and within the Transregional Collaborative Research Center TRR 173/3 - 268565370 “Spin$+$X” (Project A03), by JSPS KAKENHI (Grants No. 20H00337, No. 22K14008, No. 23H04522, and No. 24H02231), JST CREST (Grant No. JPMJCR20T1), and by Waseda University Grant for Special Research Project (Grant No. 2024C-153). R. E. was supported by a Grant-in-Aid for JSPS Fellows (Grant No. 23KJ2047) and the JSPS Overseas Challenge Program for Young Researchers FY2024. M. G. acknowledges support by the Theory of Quantum Matter Unit of the Okinawa Institute of Science and Technology Graduate University (OIST), and by the Scientific Computing section of the Research Support Division at OIST for providing the HPC resources. A part of numerical simulations was carried out at the Supercomputer Center, Institute of Solid State Physics, University of Tokyo. The work of A. L. C. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0021221.
Methods
Spin-wave expansion for collinear isotropic magnets
[…]
Gratifyingly, the above observation of a spin conservation does not have to be put into the spin-wave expansion by hand but rather emerges naturally when carrying out the expansion, as we will show in the following. We exclude the trivial case of ferromagnets. Being interested in the low-energy physics, we assume without loss of generality that there are only two sublattices, called A and B and carrying spins of length $S_{\text{A}}$ and $S_{\text{B}}$, such that there are no optical magnon branches. By means of the Holstein-Primakoff transformation [25],
$S^{z}_{{\boldsymbol{r}}\in\text{A}}=S_{\text{A}}-a^{\dagger}_{{\boldsymbol{r}}}a_{{\boldsymbol{r}}}\quad$ $\text{and}\quad S^{z}_{{\boldsymbol{r}}\in\text{B}}=-S_{\text{B}}+b^{\dagger}_{{\boldsymbol{r}}}b_{{\boldsymbol{r}}}$ (13)
$S^{-}_{{\boldsymbol{r}}\in\text{A}}=a^{\dagger}_{{\boldsymbol{r}}}\sqrt{2S_{\text{A}}-a^{\dagger}_{{\boldsymbol{r}}}a_{{\boldsymbol{r}}}}\quad$ $\text{and}\quad S^{-}_{{\boldsymbol{r}}\in\text{B}}=\sqrt{2S_{\text{B}}-b^{\dagger}_{{\boldsymbol{r}}}b_{{\boldsymbol{r}}}}\,b_{{\boldsymbol{r}}}$ (14)
we map the spin operators onto sublattice bosons with annihilators $a_{{\boldsymbol{r}}}$ and $b_{{\boldsymbol{r}}}$ and creators $a^{\dagger}_{{\boldsymbol{r}}}$ and $b^{\dagger}_{{\boldsymbol{r}}}$. By expanding the square root in Eq. (14)—assuming ground state expectation values of $a^{\dagger}_{{\boldsymbol{r}}}a_{{\boldsymbol{r}}}$ and $b^{\dagger}_{{\boldsymbol{r}}}b_{{\boldsymbol{r}}}$ small compared to $2S_{\text{A/B}}$—we obtain an infinite expansion $H=H_{0}+H_{2}+H_{4}+H_{6}+\ldots$, where the subscript denotes the power in bosonic operators. There are no terms with an odd power of bosons because of the residual SO$(2)$ symmetry around the collinear axis. (More generally, a less stringent requirement for the absence of odd terms in the spin-wave expansion of collinear magnets is a $C_{2}$ symmetry around the collinear axis [11, 3, 48]. In this case, however, the even terms do not have to conserve spin.) $H_{0}$ is the classical ground state energy that is of no further interest. $H_{2}$ encodes the harmonic theory of noninteracting magnons, and $H_{4}$ and beyond captures magnon-magnon interactions. In momentum space, the harmonic theory is block diagonal,
$H_{2}=\frac{1}{2}\sum_{{\boldsymbol{k}}\in\text{BZ}}\left(a^{\dagger}_{{\boldsymbol{k}}},b_{-{\boldsymbol{k}}},b^{\dagger}_{{\boldsymbol{k}}},a_{-{\boldsymbol{k}}}\right)\begin{pmatrix}h_{+}({\boldsymbol{k}})&0\\ 0&h_{-}({\boldsymbol{k}})\end{pmatrix}\begin{pmatrix}a_{{\boldsymbol{k}}}\\ b^{\dagger}_{-{\boldsymbol{k}}}\\ b_{{\boldsymbol{k}}}\\ a^{\dagger}_{-{\boldsymbol{k}}}\end{pmatrix}+\delta H_{0},$ (15)
where the two blocks correspond to opposite spin quantum numbers ±$\pm 1$. Here, $\delta H_{0}$ is a number providing a correction to the ground state energy, and the two-by-two Hamilton matrix blocks read
$h_{\pm}({\boldsymbol{k}})=\begin{pmatrix}A_{\pm}({\boldsymbol{k}})&B_{\pm}({\boldsymbol{k}})\\ B^{\ast}_{\pm}({\boldsymbol{k}})&C_{\pm}({\boldsymbol{k}})\end{pmatrix},$ (16)
where the entries $A_{\pm}({\boldsymbol{k}})$, $B_{\pm}({\boldsymbol{k}})$, and $C_{\pm}({\boldsymbol{k}})$ depend on the specifics of the system. After a Bogoliubov transformation, $H_{2}$ becomes diagonal,
$H_{2}=\delta H^{\prime}_{0}+\sum_{{\boldsymbol{k}}\in\text{BZ}}\left(\epsilon_{+}({\boldsymbol{k}})\alpha^{\dagger}_{{\boldsymbol{k}}}\alpha_{{\boldsymbol{k}}}+\epsilon_{-}({\boldsymbol{k}})\beta^{\dagger}_{{\boldsymbol{k}}}\beta_{{\boldsymbol{k}}}\right),$ (17)
where $\epsilon_{\pm}({\boldsymbol{k}})$ is the harmonic magnon energy of the spin-$\uparrow$ and spin-$\downarrow$ magnons, created by $\alpha^{\dagger}_{{\boldsymbol{k}}}$ and $\beta^{\dagger}_{{\boldsymbol{k}}}$, respectively. These bosons are the eigenmodes of the system when interactions are neglected. $\delta H^{\prime}_{0}$ is the full correction to the classical ground state energy due to harmonic quantum fluctuations.
[…]
In terms of the bosons diagonalizing the harmonic theory, $\alpha$ and $\beta$, the leading magnon-magnon interactions are given by
[…]
where we have suppressed momentum labels. The full expression in given in the Supplementary Note 2. Notably, none of the terms in Eq. (18) change the total spin—the total spin of the destroyed bosons equals that of those created. While Ref. [23] suggests the presence of terms like $\alpha^{\dagger}\alpha^{\dagger}\alpha^{\dagger}\beta$, such combinations do not actually appear. The same observation holds for $H_{6}$ and higher-order terms and is a consequence of spin conservation as discussed above.
To explore the effects of interactions in altermagnets, we set $S_{\text{A}}=S_{\text{B}}=S$ and follow the general strategy for many-body perturbation theory, e.g., see Refs. [8, 55, 56]. Formally, the small perturbative parameter of the spin-wave expansion is $1/S$, where $H_{n}$ is of order $S^{2-n/2}$. In the usual sense of many-body perturbation theory, we consider $H_{2}$ the unperturbed Hamiltonian piece and $V=V^{(1/S)}+V^{(1/S^{2})}$ the perturbation. The order-$1/S$ perturbation $V^{(1/S)}$ contains $H_{4}$ and the order-$1/S^{2}$ perturbation $V^{(1/S^{2})}$ contains $H_{6}$. For order-$1/S^{2}$ accuracy, we treat $V^{(1/S)}$ up to second-order perturbation theory and $V^{(1/S^{2})}$ up to first-order perturbation theory.
The interaction-corrected magnon dispersion $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})$ is obtained from the poles of the retarded Green’s function, $G_{\pm}({\boldsymbol{k}},\omega)=[\omega+\mathrm{i}0^{+}-\epsilon_{\pm}({\boldsymbol{k}})-\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\omega)-\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\omega)]^{-1}$, where the relevant zero-temperature single-magnon self-energies, $\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\omega)$ and $\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\omega)$, are given in the Supplementary Note 2. Within an on-shell calculation, the corrected spectrum is found to read
$\tilde{\epsilon}_{\pm}({\boldsymbol{k}})=\epsilon_{\pm}({\boldsymbol{k}})+\delta\epsilon^{(1/S)}_{\pm}({\boldsymbol{k}})+\delta\epsilon^{(1/S^{2})}_{\pm}({\boldsymbol{k}}),$ (19)
where $\delta\epsilon^{(1/S)}_{\pm}({\boldsymbol{k}})=\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\epsilon_{\pm}({\boldsymbol{k}}))$ and $\delta\epsilon^{(1/S^{2})}_{\pm}({\boldsymbol{k}})=\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\epsilon_{\pm}({\boldsymbol{k}}))$. At order $1/S$, there are only real corrections: $\delta\epsilon^{(1/S)}_{\pm}({\boldsymbol{k}})\in\mathbb{R}$. The leading contribution to the magnon lifetime $\tau_{\pm}({\boldsymbol{k}})=-1/\text{Im}[\delta\epsilon^{(1/S^{2})}_{\pm}({\boldsymbol{k}})]$ appears at order $1/S^{2}$ and is associated with the diagram in Fig. 3b. The corresponding decay rate of magnons, $\Gamma_{\pm}({\boldsymbol{k}})=\tau^{-1}_{\pm}({\boldsymbol{k}})$, is given in Eq. (9). It represents the half width at half maximum of the Lorentzian quasiparticle peak. We have explicitly verified that the real corrections to the spectrum do not gap out the Goldstone modes, that is, both $\delta\epsilon^{(1/S)}_{\pm}({\boldsymbol{k}})\to 0$ and $\mathrm{Re}\delta\epsilon^{(1/S^{2})}_{\pm}({\boldsymbol{k}})\to 0$ as $k\to 0$.
For the off-shell calculation presented in Fig. 5b, we have plotted the single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)=-\text{Im}[G_{+}({\boldsymbol{k}},\omega)+G_{-}({\boldsymbol{k}},\omega)]/\pi$. The resulting spectrum is no longer consistent within $1/S$ and captures some nonperturbative effects, such as non-Lorenzian lineshapes. This inconcsistency in $1/S$ generates a problem for the Goldstone mode that gets pushed to negative energies. Numerically, we found it necessary to suppress one particular family of self-energy diagrams to avoid this problem; this neglection has almost no visible effects on the spectrum away from the Brillouin zone origin as the magnitude of these self-energies is tiny (see Supplementary Note 2 for further details).
Linear spin-waves in checkerboard d-wave altermagnet
Applying spin-wave theory to the spin model in Eq. (8), results in the following noninteracting magnon dispersion
$\epsilon_{\pm}({\boldsymbol{k}})=S\left(\sqrt{A({\boldsymbol{k}})^{2}-B({\boldsymbol{k}})^{2}}\pm\Delta({\boldsymbol{k}})\right),$ (20)
where
$A({\boldsymbol{k}})=4J_{1}-4J_{2}+2J_{2}\left(\cos(k_{x}+k_{y})+\cos(k_{x}-k_{y})\right)$ (21)
and
$B({\boldsymbol{k}})=2J_{1}\left(\cos k_{x}+\cos k_{y}\right).$ (22)
The d-wave splitting is given by $\epsilon_{+}({\boldsymbol{k}})-\epsilon_{-}({\boldsymbol{k}})=2S\Delta({\boldsymbol{k}})$, where
$\Delta({\boldsymbol{k}})=-2\Delta\left(\cos(k_{x}+k_{y})-\cos(k_{x}-k_{y})\right).$ (23)
In the long-wavelength limit, the dispersion is expanded as follows
$\frac{\epsilon_{\pm}({\boldsymbol{k}})}{S}\approx v|{\boldsymbol{k}}|\pm 2\Delta\sin(2\varphi_{{\boldsymbol{k}}})|{\boldsymbol{k}}|^{2}+\left(\kappa+\kappa_{4\varphi}\cos(4\varphi_{{\boldsymbol{k}}})\right)|{\boldsymbol{k}}|^{3},$ (24)
where the nodal lines of the altermagnetic splitting are rotated by $\pi/4$ compared to Eq. (3), and we have explicitly included in $\kappa_{4\varphi}$ the angular dependence of the cubic term. The expansion coefficients are related to $J_{1}$ and $J_{2}$ by the following relations:
$v$ $=2\sqrt{2J_{1}(J_{1}-2J_{2})},$ (25)
$\kappa$ $=\frac{-3J_{1}^{2}+4J_{1}J_{2}+8J_{2}^{2}}{8\sqrt{2J_{1}(J_{1}-2J_{2})}},$ (26)
$\kappa_{4\varphi}$ $=\frac{J_{1}(J_{1}+4J_{2})}{24\sqrt{2J_{1}(J_{1}-2J_{2})}}.$ (27)
Throughout the text, we set $S=1/2$, assume units of meV, and explore two parameter sets:
The first set of parameters is used in Fig. 3, Fig. 4a,b, and Fig. 5: $J_{1}=1$ and $J_{2}=-0.5$, such that $v=4$, $\kappa=-3/16<0$, and $\kappa_{4\varphi}=-1/48$, realizing the negatively curved case with forbidden class-(ii) type decays. According to Eq. (5), we find $\Delta^{\star}_{\text{d-wave}}=3/(4\sqrt{2})\approx 0.53$.
The second set of parameters is used in Fig. 4c: $J_{1}=1/10$ and $J_{2}=-1$, such that $v=\sqrt{42}/5$, $\kappa=757/(80\sqrt{42})>0$, and $\kappa_{4\varphi}=-13/(80\sqrt{42})$, realizing the positively curved case supporting class-(ii) type decays. According to Eq. (7), we find $k^{\star}_{\text{d-wave}}=640/757\times\sqrt{7/6}\approx 0.913\Delta$.
For both parameter sets, $|\kappa_{4\varphi}|\ll|\kappa|$ and we expect that the results obtained from Eq. (3), which neglected $\kappa_{4\varphi}$, provide a reasonable expectation.
Matrix product operator based time evolution
We perform nonperturbative quantum simulations to obtain dynamical correlations using infinite matrix-product states [57, 58] and a time evolution based on matrix product operators [28]. See also Ref. [59, 60] for more details on the method. MPS provide an efficient representation of quantum wave functions in one spatial dimension, but have been applied widely to two-dimensional systems by wrapping the lattice onto a cylinder and winding around the one-dimensional tensor train of the MPS and MPO. The finite circumference of the cylinder discretizes reciprocal momentum as $k_{y}=2\pi/L_{y}$. On the other hand, reciprocal momenta $k_{x}$ parallel to the cylinder axis are very dense, resulting in lines of accessible momenta in reciprocal space. We illustrate both lattice geometries, their periodic boundary around the cylinder, and the corresponding Brillouin zone in Fig. 6.
Figure 6: Cylinder geometries and accessible momenta in reciprocal space in nonperturbative DMRG$+$tMPO simulations. (a) Regular square lattice with a circumference of $L_{y}=4$ unit cells. (b) $\pi/4$-rotated lattice with $L_{y}=3$. (c) Brillouin zone and accessible momenta cuts of regular (solid) and rotated lattice (dashed). The red line illustrates the Brillouin zone path used in Fig. 5a. The segment from $(-\frac{\pi}{2},\frac{\pi}{2})$ to $(-\pi,\pi)$ is taken from the path $(\frac{\pi}{2},\frac{\pi}{2})$ to $(\pi,\pi)$. Although these paths are not related by symmetry and have opposite altermagnetic splitting, the transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ is identical.
First, we obtain the ground state in MPS form using infinite DMRG [26, 57, 27, 58] and a bond dimension of $\chi=400$. Here, a small XXZ anisotropy $\lambda\sum_{\langle{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}\rangle}S^{z}_{{\boldsymbol{r}}}S^{z}_{{\boldsymbol{r}}^{\prime}}$ of $\lambda=0.02$ improves accuracy and convergence by gapping out the Nambu-Goldstone mode, and ensuring a finite local moment of $\langle S^{z}_{{\boldsymbol{r}}_{i}}\rangle\approx\pm 0.28$. After applying a local spin operator $S^{\gamma=\{x,y,z\}}_{{\boldsymbol{r}}_{\text{center}}}$ in the middle of the cylinder, we perform the iterative application of the time-evolution MPO [28] with a time step of $dt=0.125/7\approx 0.01786$, while spin-spin correlations are measured each $\Delta t=0.125$ up to a maximum time of $t_{\text{max}}=25$. We allow $\chi$ to grow adaptively throughout the time-evolution with a cap set to $\chi_{\text{max}}=600$ in order to limit required computational resources.
The objective is to compute the dynamical spin-structure factor $\mathcal{S}({\boldsymbol{k}},\omega)$ defined as,
$\mathcal{S}_{\gamma\gamma}(\bm{k},\omega)=N\int\mathrm{d}t~{}\mathrm{e}^{\mathrm{i}\omega t}\sum_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}\mathrm{e}^{\mathrm{i}{\boldsymbol{k}}\cdot({\boldsymbol{r}}-{\boldsymbol{r}^{\prime}})}~{}C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)~{},$ (28)
where $\gamma=\{x,y,z\}$ and $C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)=\langle S^{\gamma}_{{\boldsymbol{r}}}(t)S^{\gamma}_{{\boldsymbol{r}}^{\prime}}(0)\rangle$ are the dynamical spin-spin correlations. The protocol is as follows: (i) $C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)$ is Fourier transformed in space providing $C^{\gamma\gamma}({\boldsymbol{k}},t)$, (ii) $C^{\gamma\gamma}({\boldsymbol{k}},t)$ gets extended in time using linear predictive coding [61, 62, 63, 64], and convoluted with a Gaussian of width $\sigma_{t}\approx 11.16$ to suppress ringing from the box function of the finite time window, and (iii) in doing a Fourier transform in time and coherent summations over the two sublattices of the magnetic unit cell, we obtain $\mathcal{S}_{\gamma\gamma}({\boldsymbol{k}},\omega)$ as in Eq. (28). The convolution with a Gaussian in (ii) results in a Gaussian broadening in $\mathcal{S}_{\gamma\gamma}(\bm{k},\omega)$ of width $\sigma_{\omega}\approx 0.0896$. The final $\mathcal{S}_{\perp}(\bm{k},\omega)=\mathcal{S}_{xx}(\bm{k},\omega)+\mathcal{S}_{yy}(\bm{k},\omega)$ is plotted in Fig. 5a.
Our simulations complement those of Ref. [23]. Here, we remark on the difference to the approach taken in Ref. [23] and its implications. We compute the ground state on an infinite cylinder and perform the time-evolution on a finite cylinder segment embedded in an infinite cylinder. This is to be contrasted to the finite length, $L_{x}=80$, cylinder in Ref. [23]. Therefore, we are less affected by finite-size effects. In turn, we add a finite XXZ anisotropy, $\lambda=0.02$, to control magnetization, convergence and computational resources, while retaining interpretability of the gap. For comparison, Ref. [23] has a spin gap of $\varepsilon\approx 0.5$, while at $\lambda=0.02$ we obtain $\varepsilon\approx 0.23$. Furthermore, the ground state of the isotropic, $\lambda=0$, Heisenberg model on a finite-width cylinder does not break $\mathrm{SO}(3)$ spontaneously and develops a paramagnetic, singlet-like ground state instead. As a result, $\mathcal{S}_{\mu\mu}({\boldsymbol{k}},\omega)$ with $\mu=x,y,z$ is isotropic and picks up single-magnon and two-magnon contributions. In our case, an XXZ anisotropy of $\lambda=0.02$ ensures finite sublattice magnetization, which enables us to separate the transverse $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ from the longitudinal $\mathcal{S}_{zz}({\boldsymbol{k}},\omega)$, the latter of which we are presenting in Supplementary Note 5. However, as $\mathcal{S}_{zz}({\boldsymbol{k}},\omega)$ is proportional to the two-magnon sector rather than single particles and therefore does not directly reflect quasiparticle damping, we propose measurement of $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ in experiments.
Magnon damping estimates for $\kappa$-Cl-like altermagnets
Figure 7: Spontaneous magnon decay in $\kappa$-Cl-like organic altermagnets. (a,b) Effective spin model of $\kappa$-Cl on the square lattice derived in Ref. [21]. Magenta and cyan arrows with ellipsoids respectively denote the collinear ground state $(\pi,\pi)$ Néel order. Directions of the ellipsoids correspond to the molecular orientations, which is the origin of altermagnetism in this material. White and green bonds in (a) indicate the nearest-neighbor and the next-nearest-neighbor exchange interactions $J$ and $J^{\prime}$, respectively. In (b), due to the differing molecular orientations between the two sublattices, the third-nearest neighbor exchange interactions along the bonds parallel and perpendicular to the orientation direction are non-equivalent. Here, only the latter is denoted as $K$ and the former is neglected. (c) Renormalized magnon spectrum (lines with dot markers) compared to the linear magnon spectrum (solid lines) for $J=80\,$meV, $J^{\prime}=20\,$meV, $S=1/2$ and $K=20\,$meV. The value of $K$ sets the magnitude of the altermagnetic splitting and is approximately ten times larger than what is predicted for $\kappa$-Cl [21]. (d) Magnon damping $\Gamma_{\pm}({\boldsymbol{k}})$ along momentum directions with maximal band splitting.
$\kappa$-Cl is an effectively two-dimensional d-wave altermagnet, whose low-energy spin model was derived in Ref. [21]:
$H=\sum_{\expectationvalue{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}}J{\boldsymbol{S}}_{{\boldsymbol{r}}}\cdot{\boldsymbol{S}}_{{\boldsymbol{r}}^{\prime}}+\sum_{\expectationvalue{\expectationvalue{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}}}J^{\prime}{\boldsymbol{S}}_{{\boldsymbol{r}}}\cdot{\boldsymbol{S}}_{{\boldsymbol{r}}^{\prime}}+\sum_{\expectationvalue{\expectationvalue{\expectationvalue{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}}}}K{\boldsymbol{S}}_{\boldsymbol{r}}\cdot{\boldsymbol{S}}_{{\boldsymbol{r}}^{\prime}}.$ (29)
Here, the local spin operators ${\boldsymbol{S}}_{{\boldsymbol{r}}}$ of length $S=1/2$ are associated with molecular dimers, which are indicated by ellipsoids in Fig. 7a. $J$ and $J^{\prime}$ are the first and second-nearest-neighbor exchange interaction, respectively. $K$ is a third-nearest neighbor interaction along those third-neighbor bonds that align with the short axis of the dimers, see Fig. 7b. This third-nearest neighbor coupling causes the altermagnetic splitting in contrast to the model in Eq. (8), where it derives from second-neighbor coupling. As a result, the magnon splitting is maximal at the midpoint of the line from $(0,0)$ to $(\pi,0)$ or to $(0,\pi)$, see Fig. 7c, but it is zero at the entire boundary of the Brillouin zone. By carrying out a nonlinear spin-wave analysis identical to the one outlined above, we find that for the parameters predicted in Ref. [21], that is, $J=80\,$meV, $J^{\prime}=20\,$meV, and $K=2\,$meV, the maximal magnon damping is smaller than $1\,\mu$eV. However, if $K$ is increased by a factor of ten, we obtain a maximal damping of $\Gamma\sim 6\,$meV, as shown in Fig. 7d.
[…]
$\epsilon^{\kappa\textrm{-Cl}}_{\pm}({\boldsymbol{k}})$ $=\nu\sqrt{1+(\nu_{2\varphi}|\nu_{2\varphi}|/\nu^{2})\sin(2\varphi_{\boldsymbol{k}})}|{\boldsymbol{k}}|$
$\quad+\frac{\kappa+\kappa_{2\varphi}\sin(2\varphi_{\boldsymbol{k}})+\kappa_{4\varphi}\cos(4\varphi_{\boldsymbol{k}})}{\sqrt{1+(\nu_{2\varphi}|\nu_{2\varphi}|/\nu^{2})\sin(2\varphi_{\boldsymbol{k}})}}|{\boldsymbol{k}}|^{3}$
±$\quad\pm\Delta\cos(2\varphi_{\boldsymbol{k}})|{\boldsymbol{k}}|^{2}.$ (30)
We find that $|\nu_{2\varphi}|\ll\nu$ and $|\kappa_{2\varphi}|$, $|\kappa_{4\varphi}|\ll|\kappa|$ are satisfied. The angular dependence of the velocity $\nu_{2\varphi}$ and the cubic nonlinearity $\kappa_{2\varphi}$ stems from the next-nearest-neighbor exchange interaction $J^{\prime}$, which breaks the four-fold rotation symmetry of the square lattice. Since the crystal structure of $\kappa$-Cl belongs to the crystallographic Laue group $mmm$, which does not preserve four-fold rotational symmetry, both $\nu_{2\varphi}$ and $\kappa_{2\varphi}$ originate purely from lattice-geometric factors. We emphasize that a finite $\nu_{2\varphi}$ does not disrupt the scaling theory of $\Gamma$ in the long-wavelength limit.
We note that planar altermagnets belong to one of the following four crystallographic Laue groups: $mmm$, $4/m$, $4/mmm$, and $6/mmm$ [19]. Those belonging to $mmm$ are the only with nonzero angular corrections to the velocity, $\nu_{2\varphi}\neq 0$. In contrast, angular corrections to $\kappa$, namely $\kappa_{4\varphi}$ in d-wave and g-wave, and $\kappa_{6\varphi}$ in i-wave, are generically present. They have a qualitative effect on magnon damping if they modulate the sign of the cubic nonlinearity with direction.